# Industrial Deposition of Wavelength–Shifting Films for Liquid Argon Photon Detection Systems

### v1

@@ Abstract
The Deep Underground Neutrino Experiment (DUNE) Phase-II Far Detector is considering an unprecedentedly large-area ($\sim\!$2000 m2) photon detection system to achieve a target mean light yield of 180 PE/MeV. Meeting this requirement demands scalable, cost-effective, and high-quality wavelength-shifter (WLS) coatings capable of converting 127 nm liquid-argon scintillation light into visible photons with high efficiency. We report on the successful realization of an industrial physical vapor deposition (PVD) process for p-terphenyl (pTP) coatings, adapted from vacuum deposition techniques developed for OLED display manufacturing, to produce uniform WLS layers on large-area inorganic substrates, a task traditionally challenged by adhesion and uniformity issues at organic–inorganic interfaces. Surface characterization by profilometry and spectroscopic measurements confirms sub-10% thickness uniformity for 1-2 µm films and emission spectra consistent with high-quality pTP reference samples. The industrial process demonstrates reproducibility, scalability, and significantly reduced production time compared to laboratory-based methods, while maintaining the optical quality required for large-scale liquid-argon time-projection chamber (LArTPC) photon detection systems. These results establish a viable pathway for mass production of high-performance pTP coatings for DUNE FD3 and future neutrino experiments.
@@ 1 Introduction
In the DUNE Phase-I Far Detectors 1 and 2 (FD1/2) [2, 6], the baseline light-collection technology is the X-Arapuca light-trap design [8, 13]. This concept combines a primary wavelength-shifting (WLS) layer, an optical trapping mechanism with a secondary WLS layer, and silicon photomultiplier (SiPM) readout to achieve high photon-collection efficiency. Figure 1a illustrates the working principle of the X-Arapuca cell. In this configuration, the primary WLS is deposited directly onto a substrate facing the LAr volume, where it converts the 127 nm VUV scintillation photons into near-UV photons suitable for subsequent wavelength shift, light trapping, and detection.
For DUNE Phase-II Far Detector [1], the PDS design remains under active optimization, and several alternative light-trap geometries are being investigated. Preliminary optical simulations and prototype studies indicate that comparable photon-collection performance may be achievable even without a dichroic interface, motivating parallel development of simplified single-layer WLS coatings. The present work focuses on realizing such a coating through industrial-scale deposition of p-terphenyl (pTP) films on large-area substrates, evaluating their optical and mechanical performance as a potential solution for future DUNE Far Detector modules.
@@ 2.2 Deposition procedure for pTP coatings
Three deposition campaigns were carried out as the process was optimized. The first campaign used a generic pTP powder on small witness samples of B33, quartz, and sapphire without plasma treatment, and was employed solely to explore basic PVD operating conditions. A second campaign used full-size 143.75 mm $\times$ 143.75 mm substrates (10 sapphire, 15 B33, and 10 quartz) coated to a nominal thickness of 1.5 µm. For each substrate type, approximately half of the samples received plasma surface treatment and the remainder were processed without it, still using generic pTP. This campaign informed the choice of plasma treatment, thickness range, and chamber conditions for the final recipe.
For the third campaign, the source material was upgraded to high-purity pTP from Thermo Fisher Scientific, and only B33 substrates were coated (15 samples at 1 µm thickness), all with plasma surface treatment applied. The change from generic to branded high-purity pTP led to noticeably cleaner emission spectra with a more clearly defined peak and higher apparent light yield. Unless otherwise stated, the B33 emission and thickness results presented in the analysis in Sec. 3 focus on these best-controlled samples: high-purity pTP on B33 from the third campaign, and plasma-treated quartz and sapphire from the second campaign.
@@ 2.5 Tested substrates
As summarized in Sec. 2.2, pTP coatings were deposited on borosilicate (B33) glass, quartz (fused silica), and sapphire substrates as part of the multi-campaign optimization program. Here we briefly summarize the relevant material properties; a comparison is given in Table 1.
@@ 3.1 Emission spectrum
Figure 6: Emission spectra measured from pTP-coated samples under UV excitation. The spectra are area normalized.
Emission spectra of pTP-coated samples were acquired with long integration times to resolve the near-UV emission band; a representative set is shown in Fig. 6. All emission spectra were measured using monochromatic excitation at 266 nm, selected by the monochromator from the deuterium lamp output. As a result, the spectral shape of the excitation source does not contribute to the measured emission spectra, and no unfolding of the lamp spectrum is required.
The measured emission profiles are in agreement with the canonical pTP spectrum (Fig. 1b), both in peak position (approximately 350 nm) and overall spectral shape. Spectra obtained from independent reference samples produced at Fermilab and Stony Brook University (SBU) exhibit the same features. The Fermilab reference sample consists of Thermo Fisher pTP deposited by PVD onto a B33 glass substrate, while the SBU reference sample uses Thermo Fisher pTP deposited by a solvent-based spray process onto an acrylic substrate. The close agreement among these spectra, despite differences in substrate and deposition technique, confirms that the industrially deposited films preserve the intrinsic optical behavior of pTP.
Peak variation among all measured samples was less than 5 nm, indicating strong reproducibility across substrates and deposition batches. This consistency demonstrates that neither substrate composition nor plasma pre-treatment measurably alters the photophysical response of pTP under the conditions studied here.
During earlier process-development runs using a generic pTP powder (see Fig. 7), we observed less well-defined emission features and lower apparent light yield, indicating that control of the pTP source purity is an important factor in achieving reliable, high-quality, and reproducible emission spectra. This observation is consistent with the purity requirements discussed in Sec. 1.
@@ 3.2 Thermal and cryogenic stress response
Samples underwent multiple liquid-nitrogen cooling and warm-up cycles, after which no macroscopic delamination, peeling, or cracking was observed. Minor visual changes were noted on quartz and sapphire substrates during warm-up, possibly associated with surface condensation or transient moisture effects rather than degradation of the coating itself. Follow-up emission-spectrum measurements before and after thermal cycling are in progress to quantify any potential spectral shifts or changes in intensity. Initial observations, illustrated in Fig. 7, indicate that the emission profile remains stable within experimental uncertainty, both in spectral shape and light yield, suggesting that the industrial pTP coatings maintain optical integrity and are compatible with cryogenic operation in LArTPC photon detection systems.
@@ 3.3 Coating thickness and uniformity
Film thickness was measured by profilometry using a shallow step created by masking prior to deposition, so that the stylus scanned across a boundary between bare substrate and coated region. Multiple positions per sample were scanned to reduce bias from substrate non-flatness. As in Sec. 3.1, the B33 results shown here correspond to the optimized third deposition campaign using high-purity Thermo Fisher pTP, while the quartz and sapphire results are drawn from the plasma-treated subset of the second campaign (see Sec. 2.2).
@@
Figure 8: Profilometer results for representative B33, sapphire, and quartz substrates. Left: thickness vs. lateral scan distance, where the gray band indicates the lateral range used for thickness statistics; right: thickness distribution over 0.5–2.5 mm from the step edge.
Figure 8 illustrates typical scans. For sapphire, an apparent slope in the thickness vs. lateral distance is consistent with substrate bow rather than coating non-uniformity, complicating direct cross-sample comparisons. To standardize the metric, we report the standard deviation of thickness within the 0.5–2.5 mm interval. For each sample, six scans at different edge positions were taken; the replicate least affected by substrate flatness was selected for the summary in Table 2.
All coatings reached the target thickness range of 1–2 µm with less than 10% variation across the scanned regions. B33 glass consistently yielded the lowest standard deviations (40–60 nm), demonstrating excellent uniformity; quartz followed with <100<100 nm variation, and sapphire exhibited larger spread due to substrate bow and differing surface energetics. These results mark a significant improvement over early laboratory-scale coatings, which exhibited substantially larger thickness variations, indicating that the optimized industrial process achieves reliable large-area uniformity suitable for mass production. Such consistency across different substrates confirms the scalability and readiness of this approach for DUNE Phase-II Far Detector deployment from the coating-production standpoint.
Table 2: Representative film thickness statistics per sample and substrate. The replicate with minimal substrate-flatness bias (see text) is reported for each sample.
@@
Sample number Substrate Mean thickness (nm) Std. dev. (nm)
S2 Sapphire 1460.6 118.1
S21 Quartz 2179.9 77.9
S29 Quartz 1372.9 99.4
S35 Sapphire 926.5 188.0
S45 B33 940.2 41.5
S46 B33 828.1 39.6
@@ 4 Conclusions
This work represents, to our knowledge, the first successful demonstration of industrial-scale physical vapor deposition of $p$-terphenyl coatings on multiple optical substrates, including borosilicate glass, quartz, and sapphire, with comprehensive surface characterization confirming that the films are smooth, uniform, and adherent, and meet the optical and mechanical quality requirements for photon detection systems in liquid argon time projection chambers such as DUNE Phase-II Far Detector. The industrial process yields consistent 1–2 µm coatings with sub-10% uniformity and has proven cost-effective and highly scalable; the demonstrated throughput indicates that the full 2000 m2 coating required for DUNE could be completed within approximately one year of production.
Beyond demonstrating large-scale manufacturability, this work addresses a long-standing materials challenge: achieving uniform, adherent organic coatings on inorganic substrates. The success of the industrial deposition process shows that interface control through optimized plasma and ion treatments can effectively overcome adhesion and uniformity issues that have traditionally limited such systems. This capability validates the robustness of the present approach and opens the path to broader applications of organic scintillators and wavelength-shifting films in cryogenic detector environments. Ongoing work focuses on further improving surface uniformity to enable next-generation light-trapper geometries with enhanced photon-collection efficiency.

### v2

@@ Abstract
The Deep Underground Neutrino Experiment (DUNE) Phase-II Far Detector is considering an approximately 2000 m2 photon detection system to achieve a target mean light yield of 180 PE/MeV. Meeting this requirement demands scalable, cost-effective, and high-quality wavelength-shifter (WLS) coatings capable of converting 127 nm liquid-argon scintillation light into visible photons with controlled and reproducible optical performance. We report on the successful realization of an industrial physical vapor deposition (PVD) process for p-terphenyl (pTP) coatings, adapted from vacuum deposition techniques developed for OLED display manufacturing, to produce uniform WLS layers on large-area inorganic substrates, a task traditionally challenged by adhesion and uniformity issues at organic–inorganic interfaces. Surface characterization by profilometry and spectroscopic measurements demonstrates edge-region thickness variation below 10% and emission spectra consistent with high-quality pTP reference samples. The industrial process demonstrates reproducibility, scalability, and significantly reduced production time compared to laboratory-based methods, while maintaining optical characteristics consistent with established pTP reference samples. These results establish a viable pathway for mass production of high-performance pTP coatings for DUNE FD3 and future neutrino experiments, from a coating manufacturing and process standpoint. Detector-level performance validation, including quantitative VUV conversion efficiency measurements at 127 nm, is identified as future work.
@@ 1 Introduction
In the DUNE Phase-I Far Detectors 1 and 2 (FD1/2) [2, 6], the baseline light-collection technology is the X-Arapuca light-trap design [14]. This concept combines a primary wavelength-shifting (WLS) layer, an optical trapping mechanism with a secondary WLS layer, and silicon photomultiplier (SiPM) readout to achieve high photon-collection efficiency. Figure 1a illustrates the working principle of the X-Arapuca cell. In this configuration, the primary WLS is deposited directly onto a substrate facing the LAr volume, where it converts the 127 nm VUV scintillation photons into near-UV photons suitable for subsequent wavelength shift, light trapping, and detection.
For DUNE Phase-II Far Detector [1], the PDS design remains under active optimization, and several alternative light-trap geometries are being investigated. Preliminary optical simulations and prototype studies indicate that comparable photon-collection performance may be achievable even without a dichroic interface, motivating parallel development of simplified single-layer WLS coatings. The present work focuses on realizing such a coating through industrial-scale deposition of p-terphenyl (pTP) films on large-area substrates, evaluating their optical and mechanical properties from a manufacturing standpoint. The scope of this work is the demonstration of a scalable industrial deposition process and characterization of the resulting films’ optical and geometric properties; detector-level performance validation, including quantitative VUV conversion efficiency at 127 nm, is identified as future work.
@@ 2.2 Deposition procedure for pTP coatings
Three deposition campaigns were carried out as the process was optimized. BOROFLOAT® 33 borosilicate glass (B33) is the baseline optical filter material currently implemented in the DUNE Phase-I photon detection system and therefore serves as the primary substrate of interest in this study. Quartz and sapphire substrates were included to evaluate the general applicability and robustness of the coating technique on alternative optical materials, as well as to explore potential suitability for other detector or photonic applications. The first campaign (Batch 1) used a generic pTP powder on small witness samples of B33, quartz, and sapphire without plasma treatment, and was employed solely to explore basic PVD operating conditions. A second campaign (Batch 2) used full-size 143.75 mm $\times$ 143.75 mm substrates (10 sapphire, 15 B33, and 10 quartz) coated to a nominal thickness of 1.5 µm. For each substrate type, approximately half of the samples received plasma surface treatment and the remainder were processed without it, still using generic pTP. This campaign informed the choice of plasma treatment, thickness range, and chamber conditions for the final recipe.
For the third campaign (Batch 3), the source material was upgraded to high-purity pTP from Thermo Fisher Scientific, and only B33 substrates were coated (15 samples at 1 µm thickness), all with plasma surface treatment applied. The change from generic to branded high-purity pTP led to noticeably cleaner emission spectra with a more clearly defined peak and higher apparent light yield. Unless otherwise stated, the B33 emission and thickness results presented in the analysis in Sec. 3 focus on these best-controlled samples: high-purity pTP on B33 from the third campaign, and plasma-treated quartz and sapphire from the second campaign.
@@
Industrial durability and Quality Control (QC)
Coating durability was evaluated following a subset of the MIL-C-48497A standard for optical coatings[20]. The pTP films passed adhesion testing (no delamination after rapid removal of applied tape) and moderate abrasion testing (no visible damage after repeated cheesecloth rubbing under pressure). These results establish baseline mechanical robustness consistent with industrial optical-coating practices.
@@ 2.5 Tested substrates
As summarized in Sec. 2.2, pTP coatings were deposited on B33 glass, quartz (fused silica), and sapphire substrates as part of the multi-campaign optimization program. Here we briefly summarize the relevant material properties; a comparison is given in Table 1.
@@ 3.1 Emission spectrum
(a) Area un-normalized spectra
(b) Area normalized spectra
Figure 6: Emission spectra measured from pTP-coated samples under UV excitation at 266 nm. (a) Un-normalized spectra under identical acquisition conditions, enabling relative fluorescence yield comparison. (b) Area-normalized spectra, facilitating spectral shape comparison across samples and substrates.
Emission spectra of pTP-coated samples were acquired with long integration times using the monochromator setup described in Sec. 2.3, with monochromatic excitation at 266 nm selected from a deuterium lamp; a representative set is shown in Fig. 6.
pTP has strong absorption in the deep-UV region, peaking near 266 nm and remaining efficient below 300 nm as shown in Fig. 1b. Therefore, 266 nm was selected as a common excitation wavelength for emission measurements. For organic wavelength shifters, once the excitation energy exceeds the lowest singlet excited-state threshold, the emission spectrum is independent of excitation wavelength (Kasha’s rule [12]). As a result, the spectral shape of the excitation source does not contribute to the measured emission spectra, and no unfolding of the lamp spectrum is required. We note that excitation at 127 nm (LAr scintillation) or 178 nm (LXe scintillation) may introduce differences in absorption depth and radiative transport. Direct validation under VUV excitation is planned in a future upgrade with the full optical path and sample operated in an argon atmosphere to enable measurements at 127 nm and 178 nm.
The measured emission profiles are in agreement with the canonical pTP spectrum (Fig. 1b), both in peak position (approximately 350 nm) and overall spectral shape. Spectra obtained from independent reference samples produced at Fermilab and Stony Brook University (SBU) exhibit the same features. The Fermilab reference sample consists of Thermo Fisher pTP deposited by PVD onto a B33 glass substrate following the DUNE Phase-I design (including a dichroic filter), while the SBU reference sample uses Thermo Fisher pTP deposited by a solvent-based spray process onto an acrylic substrate. The close agreement among these spectra, despite differences in substrate and deposition technique, confirms that the industrially deposited films preserve the intrinsic optical behavior of pTP.
Since all samples were measured under identical acquisition conditions, the un-normalized spectra in Fig. 6a provide a relative fluorescence yield comparison. The B33 samples from the third deposition campaign (high-purity Thermo Fisher pTP) show fluorescence yield broadly comparable to the reference samples, and somewhat higher than the FNAL B33 reference; the latter difference may be partly attributable to the dichroic filter present in the FNAL sample. The quartz and sapphire samples from the second campaign show somewhat different spectral shape compared to the B33 and reference samples, consistent with the use of generic pTP source material in that campaign, as discussed below.
Peak variation among the B33 samples from the Batch 3 and reference samples was less than 5 nm, indicating strong reproducibility across substrates and deposition batches. This consistency demonstrates that neither substrate composition nor plasma pre-treatment measurably alters the photophysical response of high-purity pTP under the conditions studied here.
During earlier process-development runs using a generic pTP powder (see Fig. 6b), we observed less well-defined emission features and lower apparent light yield, indicating that control of the pTP source purity is an important factor in achieving reliable, high-quality, and reproducible emission spectra. This observation is consistent with the purity requirements discussed in Sec. 1.
@@ 3.2 Thermal and cryogenic stress response
Three samples were selected for cryogenic test by immersion into liquid nitrogen (LN2) using an open-mouth dewar used for DUNE cold electronics testing. Both vertical and horizontal orientation measurements were conducted to simulate filter configurations in large detectors. For the vertical configuration, samples were placed in a basket above the LN2 surface and pre-cooled for approximately 30 minutes. The basket was then slowly lowered into the liquid over $\approx$ 10 minutes to minimize thermal shock. The samples remained fully immersed for $\approx$ 18 hours. After immersion, they were raised in stages above the liquid surface to allow gradual warming, removed once the surface temperature exceeded 0 C, dried with forced air, and subsequently placed in a $\sim$30 C dryer for 1 hour. For the horizontal configuration, the same immersion and warming procedure was followed with a longer pre-cooling time of $\approx$ 1 hour prior to immersion to further reduce thermal stress at horizontal arrangement. This test was conducted after the substrates had been warmed up for 24 hours in air after completely drying.
@@
After multiple cryogenic cycles, no macroscopic delamination, peeling, or cracking was observed. Minor visual changes were noted on quartz and sapphire substrates during warm-up, possibly associated with surface condensation or transient moisture effects rather than degradation of the coating itself. A direct comparison of the emission spectra measured before and after thermal cycling under identical, non-normalized acquisition settings for one B33 sample is shown in Fig. 7. Within experimental uncertainty, no significant change in spectral shape or emission intensity was observed. While these results are encouraging, we emphasize that this study represents a preliminary assessment based on a limited number of samples. More controlled long-duration cryogenic tests, including operation in a moisture-free environment, additional profilometric surface characterization, and studies across a larger number of tiles, are planned to further evaluate long-term robustness for large-scale LArTPC deployment.
@@ 3.3 Coating thickness and uniformity
Figure 8: Representative profilometer scans for each substrate type. Left: thickness vs. lateral scan distance, where the gray band indicates the lateral range used for thickness statistics (0.5–2.5 mm from the step edge). Right: thickness distribution over the same interval. (a, b) B33 glass (Sample S46, Batch 3); (c, d) Sapphire (Sample S35, Batch 2); (e, f) Quartz (Sample S21, Batch 2). The apparent slope in the sapphire scan (c) is consistent with substrate bow rather than coating non-uniformity. Full thickness statistics across all measured samples are reported in Table 2.
Film thickness was measured using the stylus profilometer described in Sec. 2.4, by scanning across a masked step between the uncoated and coated regions of each substrate; the central 0.5–2.5,mm interval of each scan was used for thickness statistics. For Batch 2 samples, six scans at different edge positions distributed around the substrate perimeter were taken; for Batch 3 samples, up to four corner positions were available due to the placement of the masked step. Table 2 reports statistics computed across all valid scans for every measured sample in both campaigns; samples excluded from the table and the reasons for exclusion are noted in the caption. Two complementary statistics are reported per sample: the scan-to-scan standard deviation, quantifying thickness variation between different edge positions on the same substrate, and the per-scan standard deviation (reported as a min–max range across scans), quantifying local height fluctuation within each individual 0.5–2.5 mm scan interval. We note that the stylus profilometer is limited in travel range to mm-scale scans near the masked step edge, and therefore these measurements do not constitute a full-area uniformity map across the $\times143.75 mm$ substrate.
[…]
@@
Batch Sample Substrate No. scans Mean thickness (nm) Scan-to-scan std dev (nm) Per-scan std dev range (nm)
2 S9 B33 6 1758.5 210.9 50.8 – 396.6
2 S12 B33 6 1585.6 46.6 56.3 – 108.4
2 S15 B33 6 2115.3 165.2 41.6 – 239.4
2 S20 B33 6 1526.2 136.9 54.4 – 147.8
3 S40 B33 3 987.1 72.6 56.5 – 173.3
3 S42 B33 4 990.5 18.5 50.9 – 131.8
3 S43 B33 4 1001.3 39.4 27.5 – 173.9
3 S44 B33 4 960.3 47.7 61.4 – 140.1
3 S45 B33 2 965.9 36.4 41.5 – 55.4
3 S46 B33 4 955.1 88.7 33.8 – 137.2
2 S21 Quartz 6 2124.0 160.0 77.9 – 297.0
2 S29 Quartz 6 1336.1 89.1 65.6 – 229.1
2 S2 Sapphire 6 1562.3 187.1 118.1 – 491.8
2 S35 Sapphire 6 1028.6 300.5 119.8 – 514.2
Fewer scans available due to insufficiently large masked step at some corners.
Figure 8 illustrates typical scans. For sapphire, an apparent slope in the thickness vs. lateral distance is consistent with substrate bow rather than coating non-uniformity, complicating direct cross-sample comparisons. This is also reflected in the large per-scan standard deviation range observed for sapphire samples in Table 2, where individual scans at different edge positions differ substantially due to substrate bow rather than actual coating non-uniformity. The full set of per-sample statistics across all valid scans is reported in Table 2.
All coatings reached the target thickness range of 1–2 µm. Batch 3 B33 samples show highly consistent mean thicknesses in the range 955–1001 nm across six samples, with scan-to-scan standard deviations of 18–89 nm, demonstrating good reproducibility across the batch. Batch 2 B33 samples span a broader range of 1526–2115 nm, reflecting the wider nominal thickness target of 1.5 µm used in that campaign. Quartz samples range from 1336–2124 nm, and sapphire from 1029–1562 nm. The per-scan standard deviation for B33 samples is generally well below 10% of the mean thickness across the scanned edge regions, while sapphire exhibits larger spread attributable primarily to substrate bow rather than coating non-uniformity. We note, however, that these uniformity statistics are derived from mm-scale scans near the substrate edges and do not characterize the full 143.75 mm $\times$ 143.75 mm substrate area.
Such consistency across different substrates and between batches is encouraging from a manufacturing standpoint, though full-area uniformity mapping across the bulk of each substrate remains to be demonstrated.
@@ 4 Conclusions
This work represents, to our knowledge, the first successful demonstration of industrial-scale physical vapor deposition of $p$-terphenyl coatings on multiple optical substrates, including borosilicate glass, quartz, and sapphire, with comprehensive surface characterization confirming that the films are smooth, uniform, and adherent, and meet the optical and geometric quality requirements demonstrated in this manufacturing and process study. The industrial process yields consistent 1–2 µm coatings with sub-10% uniformity and has proven cost-effective and highly scalable; the demonstrated throughput indicates that the full 2000 m2 coating required for DUNE could be completed within approximately one year of production.
Beyond demonstrating large-scale manufacturability, this work addresses a long-standing materials challenge: achieving uniform, adherent organic coatings on inorganic substrates. The success of the industrial deposition process shows that interface control through optimized plasma and ion treatments can effectively overcome adhesion and uniformity issues that have traditionally limited such systems. This capability validates the robustness of the present approach and opens the path to broader applications of organic scintillators and wavelength-shifting films in cryogenic detector environments. Ongoing work focuses on further improving surface uniformity to enable next-generation light-trapper geometries with enhanced photon-collection efficiency. Future work will also include quantitative measurements of VUV photon conversion efficiency at 127 nm and 178 nm (relevant to Xe-doped LAr scenarios), which will connect the demonstrated film quality to detector-level light-yield requirements.